The twcal package^{∗} for use with L^{A}TEX2e

Gerd Neugebauer Mainzer Str. 16 56321 Rhens

             gerd@informatik.uni-koblenz.de

1996/07/18

1 Introduction

The font twcal provides a calligraphic font. This font has been written by Thomas Weinmann. It can be found on the CTAN in the directory tex-archive/fonts/twcal. This package provides means to use this font.

2 Usage

This file can be used as a package by placing its name in the argument of \usepackage. Afterwards the font family twcal is defined. This could also have been done by providing a font definition file.

\twcal The command \twcal changes the current font family to twcal and the encoding to OT1. Usually this should be used in a TEX group only since it changes the font encoding and redefines the umlaut macro.

\texttwcal The macro \texttwcal typesets its argument in the twcal font. \twcallogo The command \twcallogo typesetz the logo for the twcal font as defined by

the author of the font.

3 Implementation

First we declare a new font family for the twcal font.

 1 \DeclareFontFamily{OT1}{twcal}{}

This font is only available in the normal shape. Here we can get the desired font by (silently) scaling the only present fonts.

∗This file has version number v1.2, dated 1996/07/18.

 2 \DeclareFontFamily{OT1}{twcal}{}
 3 \DeclareFontShape{OT1}{twcal}{m}{n}{<->s*[0.7]twcal14}{}

Next we define some macros to perform the mapping of umlaut characters to the unusual positions used in this font.

 4 \newcommand\twcal@a{\char26}
 5 \newcommand\twcal@o{\char27}
 6 \newcommand\twcal@u{\char28}
 7 \newcommand\twcal@A{\char29}
 8 \newcommand\twcal@O{\char30}
 9 \newcommand\twcal@U{\char31}

Now we define the font changing command.

\twcal The macro \twcal selects the twcal family.

10 \DeclareRobustCommand\twcal{%
11 \let\twcal@original@umlaut=\"
12 \def\"##1{%
13  \@ifundefined{twcal@##1}{%
14  \twcal@original@umlaut{##1}}{%
15  \@nameuse{twcal@##1}}}%
16 \fontfamily{twcal}%
17 \fontencoding{OT1}%
18 \selectfont}

\texttwcal The macro \texttwcal typesets its argument in the twcal font.

19 \newcommand\texttwcal[1]{\begingroup\twcal #1\endgroup}

\twcallogo Finally we provide the logo for this font as defined by the author of the font.

20 \newcommand\twcallogo{\mbox{\({
21  \lower-0.3ex\hbox{T}\kern-0.22em
22  \lower0.3ex\hbox{W}\kern-0em
23  \lower-0.3ex\hbox{C}\kern-0.24em
24  \lower0.3ex\hbox{A}\kern-0.14em
25  \lower-0.3ex\hbox{L}}\kern-0.19em
26  \lower0.1ex\hbox{1}\kern-0.14em
27  \lower-0.1ex\hbox{4}\)}}